OpenStudy (anonymous):

how much more area does a large pizza with a 12in diameter have than a small pizza with an 8 in diameter

OpenStudy (anonymous):

Area of a circle = pi * r^2, where r is the radius and pi is about 3.14

OpenStudy (anonymous):

pi * 12^2 - pi * 8^2 = pi * (144 - 64) = pi * (80) = 80 * pi

OpenStudy (anonymous):

The answer is 80 times pi.

OpenStudy (anonymous):

you need to subtract, not divide here

OpenStudy (anonymous):

same

OpenStudy (anonymous):

80 isnt one of the answer choices

OpenStudy (anonymous):

its 87, 63, 38, 13

OpenStudy (anonymous):

your question was poorly phrased

OpenStudy (dan815):

my bad i didnt notice they were diameter

OpenStudy (anonymous):

aha!

OpenStudy (anonymous):

me neither

OpenStudy (anonymous):

36 - 16 = 20pi

OpenStudy (anonymous):

what would my answer be

OpenStudy (dan815):